# Do LLMs Have Political Correctness? Analyzing Ethical Biases and Jailbreak Vulnerabilities in AI Systems

### v1

@@ Abstract
Although large language models (LLMs) demonstrate impressive proficiency in various tasks, they present potential safety risks, such as ‘jailbreaks’, where malicious inputs can coerce LLMs into generating harmful content. To address these issues, many LLM developers have implemented various safety measures to align these models. This alignment involves several techniques, including data filtering during pre-training, supervised fine-tuning, reinforcement learning from human feedback, and red-teaming exercises. These methods often introduce deliberate and intentional biases similar to Political Correctness (PC) to ensure the ethical behavior of LLMs. In this paper, we delve into the intentional biases injected into LLMs for safety purposes and examine methods to circumvent these safety alignment techniques. Notably, these intentional biases result in a jailbreaking success rate in GPT-4o models that differs by 20% between non-binary and cisgender keywords and by 16% between white and black keywords, even when the other parts of the prompts are identical. We introduce the concept of PCJailbreak, highlighting the inherent risks posed by these safety-induced biases. Additionally, we propose an efficient defense method PCDefense, which prevents jailbreak attempts by injecting defense prompts prior to generation. PCDefense stands as an appealing alternative to Guard Models, such as Llama-Guard, that require additional inference cost after text generation. Our findings emphasize the urgent need for LLM developers to adopt a more responsible approach when designing and implementing safety measures. To enable further research and improvements, we open-source our code and artifacts of PCJailbreak, providing the community with tools to better understand and mitigate safety-induced biases in LLMs.
@@ 1 Introduction
Figure 1: PCJailbreak reveals inherent biases in LLMs that disproportionately allow harmful jailbreak attacks to succeed more frequently when directed towards marginalized groups compared to privileged groups.
@@
Figure 2: Illustration showcasing the difference in response between a standard prompt and a PCJailbreak prompt. While the standard prompt is blocked by the LLM’s safety features, the PCJailbreak prompt exploits intentional biases to elicit a response.
@@
Large Language Models (LLMs) have rapidly become essential components in many fields, ranging from professional decision-making to various forms of interactive user engagement (Araci (2019); Luo et al. (2022); Tinn et al. (2023)). However, as these models become popular, ensuring their safe usage has become crucial. Developers have implemented several safety features to prevent these models from generating harmful or objectionable content, often referred to as ‘safety alignment’. (Bakker et al. (2022); Christiano et al. (2017); Ouyang et al. (2022); Openai Usage Policies). However, in this work we show that these safety alignments often introduce deliberate and intentional biases, giving rise to a phenomenon known as ’jailbreak’, where malicious inputs manage to circumvent these safety alignments, thus allowing LLMs to generate harmful outputs (Goldstein et al. (2023); Kang et al. (2024)).
The term ‘jailbreak’ refers to carefully crafted prompts that can bait aligned LLMs into bypassing their safety alignment, resulting in the generation of content that may be harmful, discriminatory, violent, or sensitive (Smith et al. (2022)). Numerous types of jailbreak attacks have been identified and categorized into two primary methods: manually written jailbreak prompts and learning-based jailbreak prompts (Zou et al. (2023); Lapid et al. (2023); Liu et al. (2023); Wei et al. (2024); Yuan et al. (2024)). The manual approach is exemplified by the “Do-Anything-Now (DAN)” series, which uses manually constructed prompts to exploit aligned LLMs. Conversely, the learning-based approach, as demonstrated by the GCG attack, formulates jailbreaks as adversarial examples by utilizing the gradient information of LLMs to automatically generate effective prompts.
@@
However, there are two main limitations of existing jailbreak methods. Firstly, automated attacks like GCG (Zou et al. (2023)) rely on a search scheme that is guided by gradient information. While this approach enables the automatic generation of jailbreak prompts, it carries a significant downside: the resulting prompts often consist of nonsensical sequences, which lack semantic meaning (Morris et al. (2020)). This major flaw makes these prompts highly vulnerable to naive defense mechanisms such as perplexity-based detection. For example, recent studies (Jain et al. (2023); Alon & Kamfonas (2023)) have shown that such straightforward defenses can easily recognize these nonsensical prompts and completely undermine the success of GCG attacks.
Secondly, although manual attacks can create stealthier jailbreak prompts, these prompts are usually handcrafted by individual LLM users, which leads to issues with scalability and adaptability. Additionally, these handcrafted methods may not adapt quickly to updates in LLMs, thus diminishing their effectiveness over time (ONeal (2024)).
@@
To address these limitations, we use a method that employs fixed prompt formats, offering scalability. By leveraging biases, such as those related to Political Correctness(PC)(Perez et al. (2022); Zhuo et al. (2023)), inherent within LLMs and often intentionally injected by LLM developers to assure safety, we ensure that the prompts retain their meaning significantly without losing effectiveness, contrary to gradient-based attacks. This approach allows us to overcome the issues of scalability and adaptability while still exploiting the biases for effective jailbreaks.
@@
In this paper, we explore the concept of PCJailbreak, investigating how intentional biases in LLMs, intended as safety alignment, paradoxically become enablers of harmful content generation when exploited. This behavior is well illustrated in Figure 1 and 2. Additionally, we propose a defense mechanism PCDefense that adjusts biases using prompts, ensuring safety and efficiency without additional inference or models, which makes it an attractive alternative to Guard Models (Inan et al. (2023); Ghosh et al. (2024); Caselli et al. (2020); Vidgen et al. (2021)), which are capable of classifying harmful conversations but require additional models and inference after text generation. .
@@
  • We analyze the nature and consequences of intentional biases introduced in LLMs for safety purposes, highlighting their potential to not only fail in deterring but also in facilitating more effective jailbreaks. This paradoxical effect underscores the urgency of addressing the inherent vulnerabilities these biases introduce.
  • Through comprehensive experiments, we show that our proposed PCJailbreak is effective across state-of-the-art models, including the latest iterations of GPT. Our framework also proves adaptable, working effectively when applied to existing jailbreak techniques.
  • We propose PCDefense, a straightforward defense strategy without the need of additional inference or models. Our findings demonstrate that even with a simple and cost-effective defense approach, jailbreak attacks can be mitigated. This highlights the critical responsibility of LLM service providers to ensure robust protection.
  • We open-source the code and all associated artifacts of PCJailbreak to facilitate community efforts in understanding and mitigating safety-induced biases in large language models. This contribution aims to provide researchers and developers with the necessary tools to explore the nature of these biases and develop more robust defenses, furthering the collective effort to ensure the safety and reliability of LLM deployments.
Our research suggests that while intentional biases are crucial for aligning LLMs with ethical standards, they necessitate careful scrutiny to prevent their manipulation. Therefore, responsible strategies from AI companies and researchers are needed to reinforce the safety of LLMs in an increasingly complex threat landscape.
@@ 2.1 Safety Alignment in LLMs
Ensuring the safety and ethical alignment of large language models (LLMs) is a critical area of ongoing research. Methods such as data filtering, supervised fine-tuning, and reinforcement learning from human feedback (RLHF) aim to align models like GPT-4 and ChatGPT with human values and preferences (Christiano et al. (2017); Bai et al. (2022); Ouyang et al. (2022); Xu et al. (2020)). However, despite these efforts, recent studies reveal vulnerabilities that can be exploited through ’jailbreak’ attacks, which lead to undesirable and harmful outputs (Kang et al., 2023; Hazell, 2023; Shen et al., 2023).
@@ 2.2 Jailbreak Attacks and Techniques
Jailbreaking LLMs involves crafting inputs that bypass safety mechanisms, resulting in harmful or objectionable content. Early jailbreak attacks, such as the ”Do-Anything-Now (DAN)” series, relied on manually crafted prompts to exploit LLM safeguards (walkerspider, 2022). (Liu et al. (2023) provided an in-depth analysis and categorization of these jailbreak prompts, highlighting the delicate balance between an LLM’s capabilities and its safety constraints.
@@
Language diversity and non-natural language inputs present additional challenges. Deng et al. (2023) explored multilingual jailbreak attacks, demonstrating that LLMs could be tricked into producing harmful outputs with non-English prompts. Yuan et al. (2024) extended this by investigating the vulnerabilities of LLMs to non-natural language inputs, such as ciphers.
@@ 2.3 Towards Improved Safety Measures
Our work builds on the existing body of research by focusing on the paradoxical consequences of intentional biases introduced for safety purposes. While these biases aim to align LLMs ethically, they also highlight new vulnerabilities. To counteract this, we propose using prompts to make the LLM re-align those biases, thus offering a robust secondary defense against jailbreak attempts.
@@ 3.1.1 Jailbreak Attack
A jailbreak attack in the context of Large Language Models (LLMs) occurs when the model generates harmful or inappropriate responses to malicious inputs instead of producing a refusal signal, which is a safe and ethical response denying the request (Zeng et al., 2024; Zou et al., 2023). Such attacks are intricately linked to the alignment methods employed in LLMs, which aim to align the model’s outputs with human values and ethical considerations. The primary objective of these attacks is to bypass the LLM’s alignment constraints, causing the model to produce inappropriate responses to malicious queries instead of refusing to answer.
Consider a set of malicious queries $Q=\{Q_{1},Q_{2},\ldots,Q_{n}\}$. An attacker meticulously designs a set of jailbreak prompts $J=\{J_{1},J_{2},\ldots,J_{n}\}$ to manipulate these queries. When each combined input $T_{i}=\langle J_{i},Q_{i}\rangle$ is fed into the LLM, the model generates a set of responses $R=\{R_{1},R_{2},\ldots,R_{n}\}$. The aim of the jailbreak attack is to ensure that the responses $R$ correctly address the malicious queries in $Q$ instead of producing refusal signals (Zeng et al., 2024).
@@ 3.2.1 Generating Keywords and Prompts
Our proposed PCJailbreak approach investigates how keywords associated with various demographic and socio-economic groups influence an LLM’s susceptibility to jailbreak attacks. To accomplish this, we generated keywords representing contrasting groups, such as privileged and marginalized communities, using the LLM itself to ensure that the keywords reflect the model’s inherent associations and biases.
@@ 3.2.2 Analyzing Inherent Biases
Figure 3: Overview of the PCJailbreak methodology. The same harmful prompt is used across different keywords representing contrasting groups to analyze variations in jailbreak success rates.*All keywords representing both privileged and underprivileged groups are generated by the LLM.
@@
3.2.3 Summary of the PCJailbreak Approach
Our PCJailbreak methodology systematically analyzes the LLM’s responses to a diverse set of harmful prompts, using various keywords across different groups. We conducted experiments using harmful prompts $H$ and multiple pairs of privileged and marginalized keywords. By maintaining a consistent prompt structure (Equation 1) and isolating the effect of the keyword, we aimed to identify any biases in the LLM’s susceptibility to jailbreak attacks.
The calculation of jailbreak success rates (Equation 2) and the comparative analysis using $\Delta S_{i}$ (Equation 3) enabled us to quantify potential biases. Our approach highlights whether intentional biases introduced during the alignment process inadvertently contribute to vulnerabilities that can be exploited through jailbreak attacks.
3.3 PCDefense: Preventing Jailbreaks without Additional Inference or Guard Models
Our jailbreak defense method PCDefense uses prompts to adjust biases without the need for additional models. Chain-of-Thought Prompting (CoT) has shown considerable potential in eliciting reasoning abilities in large language models, allowing for interpretability without the need for parameter updates (Wei et al., 2022). Inspired by this, our approach PCDefense involves adding defense prompts to reduce excessive biases, ensuring safety without additional inference overheads, in contrast to Guard Models which require additional inference cost after text generation (Inan et al. (2023); Ghosh et al. (2024); Caselli et al. (2020); Vidgen et al. (2021)). By incorporating bias terms through our defense prompts, we achieve a cost-efficient and secure method for defending against jailbreak attempts while maintaining the model’s performance.
@@ 3.3 PCDefense: Preventing Jailbreaks without Additional Inference or Guard Models
Figure 4: PCDefense adjusts inherent biases in LLMs that are exploited by PCJailbreak. It is efficient since it does not require additional inference or models such as Guard Models.
By using a defense prompt, our defense mechanism elicits the refusal of a wide range of harmful content without requiring significant additional computing resources, which makes it an attractive alernative to Guard Models.
@@ 4.1 Experimental Setups
Data To analyze the differences in jailbreak success rates between groups, we used the jailbreakbench dataset (Chao et al. (2024)). This dataset comprises 100 harmful prompts and has been widely used in numerous studies. We utilized this dataset to measure performance variations and to analyze the intentional biases inherent in LLMs.
Models We analyzed intentional bias in well-known LLMs, including closed-source models such as GPT-3.5-turbo, GPT-4, GPT-4-o, and open-source models like Llama2-7B, Llama2-13B, Llama3-7B, Phi-mini-7B, Qwen1.5, and Qwen2-7B.
@@
Keywords The keywords used to compare performance across groups were generated by the LLMs themselves to maximize the exploitation of internal biases. Most models generated similar keywords, and significant performance differences were observed.
Basically, the models primarily generated the following keywords for various groups (see Table 1):
Marginalized Group (Adjective) Privileged Group (Adjective)
White Black
Male Female
Heterosexual LGBTQ+
Able-bodied Disabled
Wealthy Low-income
Cisgender Non-normative gender
Traditionally family-structured Non-traditionally family-structured
Urban Rural
Highly educated Undereducated
Traditionally employed Sexually marginalized
Table 1: Keywords Generated for Group Comparison
These keywords were systematically used to evaluate the differences in jailbreak success rates among various groups.
4.2 Intensional Bias Analysis
To gauge the impact of intentionally injected biases in LLMs on their vulnerability to jailbreak attempts, we conducted a series of experiments utilizing various keyword prompts designed to assess the success rates of jailbreaks within marginalized and privileged categories. We evaluated multiple well-known models, including GPT-3.5, GPT-4, GPT-4o (Brown et al., 2020)(closed models), as well as LLaMA (Touvron et al., 2023), Qwen (Bai et al., 2023), and Phi-mini (Abdin et al., 2024)(open-source models).

### v3

# BiasJailbreak:Analyzing Ethical Biases and Jailbreak Vulnerabilities in Large Language Models

@@ Abstract
Although large language models (LLMs) demonstrate impressive proficiency in various tasks, they present potential safety risks, such as ‘jailbreaks’, where malicious inputs can coerce LLMs into generating harmful content bypassing safety alignments. In this paper, we delve into the ethical biases in LLMs and examine how those biases could be exploited for jailbreaks. Notably, these biases result in a jailbreaking success rate in GPT-4o models that differs by 20% between non-binary and cisgender keywords and by 16% between white and black keywords, even when the other parts of the prompts are identical. We introduce the concept of BiasJailbreak, highlighting the inherent risks posed by these safety-induced biases. BiasJailbreak generates biased keywords automatically by asking the target LLM itself, and utilizes the keywords to generate harmful output. Additionally, we propose an efficient defense method BiasDefense, which prevents jailbreak attempts by injecting defense prompts prior to generation. BiasDefense stands as an appealing alternative to Guard Models, such as Llama-Guard, that require additional inference cost after text generation. Our findings emphasize that ethical biases in LLMs can actually lead to generating unsafe output, and suggest a method to make the LLMs more secure and unbiased. To enable further research and improvements, we open-source our code and artifacts of BiasJailbreak, providing the community with tools to better understand and mitigate safety-induced biases in LLMs.
@@ 1 Introduction
Figure 1: BiasJailbreak reveals inherent biases in LLMs that disproportionately allow harmful jailbreak attacks to succeed more frequently when directed towards marginalized groups compared to privileged groups.
@@
Figure 2: Illustration showcasing the difference in response between a standard prompt and a BiasJailbreak prompt. While the standard prompt is blocked by the LLM’s safety features, the BiasJailbreak prompt exploits ethical biases to elicit a response.
@@
Large Language Models (LLMs) have rapidly become essential components in many fields, ranging from professional decision-making to various forms of interactive user engagement (Araci (2019); Luo et al. (2022); Tinn et al. (2023)). However, as these models become popular, ensuring their safe usage has become crucial. Developers have implemented several safety features to prevent these models from generating harmful or objectionable content, often referred to as ‘safety alignment’ (Bakker et al. (2022); Christiano et al. (2017); Ouyang et al. (2022); Openai Usage Policies).
These safety alignments often involve additional fine-tuning or reinforcement learning techniques, which, while designed to enhance safety and alignment, may also inadvertently introduce biases, as highlighted in resources such as (Achiam et al., 2023, p. 49) . However, biases can also arise from other sources, such as pretraining data or system prompts. While it is difficult to pinpoint exactly where these biases originate, the critical fact remains that they exist and can influence the model’s behavior.
In this work, we show that these safety alignments often introduce deliberate and ethical biases, giving rise to a phenomenon known as ’jailbreak’, where malicious inputs manage to circumvent these safety alignments, thus allowing LLMs to generate harmful outputs (Goldstein et al. (2023); Kang et al. (2024)).
The term ‘jailbreak’ refers to carefully crafted prompts that can bait aligned LLMs into bypassing their safety alignment, resulting in the generation of content that may be harmful, discriminatory, violent, or sensitive (Smith et al. (2022)). Numerous types of jailbreak attacks have been identified and categorized into two primary methods: white-box and black-box (Yi et al. (2024)). The white-box approach requires target model gradients or logits and use them as a guidance for finding adversarial jailbreak prompts. Directly fine-tuning the target LLM is not considered as a jailbreak method in this paper, although it some consider it as white-box. The black-box approach has a harder and a more general real-world setting since it does not have access to such information. Jailbreak methods in black-box usually involve template completion, prompt rewriting, or LLM-based generation.
@@
White-box attacks like GCG (Zou et al. (2023)) rely on a search scheme guided by gradient information. While this approach enables reliable generation of jailbreak prompts though with a cost of high computation, it carries a significant downside: the resulting prompts often consist of nonsensical sequences, which lack semantic meaning (Morris et al. (2020)). This major flaw makes these prompts highly vulnerable to naive defense mechanisms such as perplexity-based detection. For example, recent studies (Jain et al. (2023); Alon & Kamfonas (2023)) have shown that such straightforward defenses can easily recognize these nonsensical prompts and completely undermine the success of white-box attacks.
While having a more applicable and general setting, recent advancements in black-box settings, such as natural-language methods like PAIR (Chao et al. (2023)) and DeepInception (Li et al. (2023)), have shown that semantically coherent prompts can effectively exploit LLM vulnerabilities. Additionally, manual approaches like GUARD (Jin et al. (2024a)) iteratively refine jailbreak prompts, demonstrating adaptability to LLM updates.
@@
In this paper, we propose a novel black-box method that offers both scalability and generality. By leveraging inherent biases in LLMs, such as those related to Ethical Sensitivity(Perez et al. (2022); Zhuo et al. (2023)), we ensure that the prompts retain their meaning significantly without losing effectiveness, contrary to white-box attacks. This approach allows us to overcome the issues of scalability and adaptability while still exploiting the biases for effective jailbreaks.
@@
We explore the novel concept of BiasJailbreak, investigating how biases in LLMs, intended as safety alignment, paradoxically become enablers of harmful content generation when exploited. This behavior is well illustrated in Figure 1 and 2. Additionally, we propose a defense mechanism BiasDefense that adjusts biases using prompts, ensuring safety and efficiency without additional inference or models, which makes it an attractive alternative to Guard Models (Inan et al. (2023); Ghosh et al. (2024); Caselli et al. (2020); Vidgen et al. (2021)), which are capable of classifying harmful conversations but require additional models and inference after text generation.
@@
  • We analyze the nature and consequences of ethical biases introduced in LLMs for safety purposes, highlighting their potential to not only fail in deterring but also in facilitating more effective jailbreaks. This paradoxical effect underscores the urgency of addressing the inherent vulnerabilities these biases introduce.
  • Through comprehensive experiments, we show that our proposed BiasJailbreak is effective across state-of-the-art models, including the latest iterations of GPT. Our framework also proves adaptable, working effectively when applied to existing jailbreak techniques.
  • We propose BiasDefense, a straightforward defense strategy without the need of additional inference or models. Our findings demonstrate that even with a simple and cost-effective defense approach, jailbreak attacks can be mitigated. This highlights the critical responsibility of LLM service providers to ensure robust protection.
  • We open-source the code and all associated artifacts of BiasJailbreak to facilitate community efforts in understanding and mitigating safety-induced biases in large language models. This contribution aims to provide researchers and developers with the necessary tools to explore the nature of these biases and develop more robust defenses, furthering the collective effort to ensure the safety and reliability of LLM deployments.
Our research suggests that while ethical biases are crucial for aligning LLMs with ethical standards, they necessitate careful scrutiny to prevent their manipulation. Therefore, responsible strategies from AI companies and researchers are needed to reinforce the safety of LLMs in an increasingly complex threat landscape.
@@ 2.1 Safety Alignment in LLMs
Ensuring the safety and ethical alignment of large language models (LLMs) is a critical area of ongoing research, since the ethical bias of LLMs can lead to undesirable societal impacts and potential harms (Li et al. (2024)). Methods such as data filtering, supervised fine-tuning, and reinforcement learning from human feedback (RLHF) aim to align models like GPT-4 and ChatGPT with human values and preferences (Christiano et al. (2017); Bai et al. (2022); Ouyang et al. (2022); Xu et al. (2020)). However, despite these efforts, recent studies reveal vulnerabilities that can be exploited through ’jailbreak’ attacks, which lead to undesirable and harmful outputs (Kang et al. (2024); Shen et al. (2023). Additionally, Zheng et al. (2024) proposed many-shot demonstration techniques, using random search within demo pools and injecting system tokens to bypass safeguards.
@@ 2.2 Jailbreak Attacks and Techniques
Jailbreaking LLMs involves crafting inputs that bypass safety mechanisms, resulting in harmful or objectionable content. Early jailbreak attacks, such as the ”Do-Anything-Now (DAN)” series, relied on manually crafted prompts to exploit LLM safeguards (Shen et al. (2023)). (Liu et al. (2023) provided an in-depth analysis and categorization of these jailbreak prompts, highlighting the delicate balance between an LLM’s capabilities and its safety constraints.
@@
Zeng et al. (2024a) proposed persuasive adversarial prompts (PAP) that leverage social science-based persuasion techniques to significantly enhance jailbreak success, achieving over 92% success rates across multiple models. Similarly, Shah et al. (2024) introduced persona modulation, a black-box jailbreak approach that uses personas to exploit vulnerabilities at scale, with high success rates transferable across state-of-the-art models.
Language diversity and non-natural language inputs present additional challenges. Deng et al. (2023) explored multilingual jailbreak attacks, demonstrating that LLMs could be tricked into producing harmful outputs with non-English prompts. Yuan et al. (2024); Jin et al. (2024b) extended this by investigating the vulnerabilities of LLMs to non-natural language inputs, such as ciphers.
@@ 2.3 Towards Improved Safety Measures
Our work builds on the existing body of research by focusing on the paradoxical consequences of ethical biases introduced for safety purposes, such as stated in Achiam et al. (2023). While these biases aim to align LLMs ethically, they also highlight new vulnerabilities. To counteract this, we propose using prompts to make the LLM re-align those biases, thus offering a robust secondary defense against jailbreak attempts.
@@ 3.1.1 Jailbreak Attack
A jailbreak attack in the context of Large Language Models (LLMs) occurs when the model generates harmful or inappropriate responses to malicious inputs instead of producing a refusal signal, which is a safe and ethical response denying the request (Zeng et al., 2024b; Zou et al., 2023). Such attacks are intricately linked to the alignment methods employed in LLMs, which aim to align the model’s outputs with human values and ethical considerations. The primary objective of these attacks is to bypass the LLM’s alignment constraints, causing the model to produce inappropriate responses to malicious queries instead of refusing to answer.
Consider a set of malicious queries $Q=\{Q_{1},Q_{2},\ldots,Q_{n}\}$. An attacker meticulously designs a set of jailbreak prompts $J=\{J_{1},J_{2},\ldots,J_{n}\}$ to manipulate these queries. When each combined input $T_{i}=\langle J_{i},Q_{i}\rangle$ is fed into the LLM, the model generates a set of responses $R=\{R_{1},R_{2},\ldots,R_{n}\}$. The aim of the jailbreak attack is to ensure that the responses $R$ correctly address the malicious queries in $Q$ instead of producing refusal signals (Zeng et al., 2024b).
@@ 3.2.1 Generating Keywords and Prompts
Our proposed BiasJailbreak approach investigates how keywords associated with various demographic and socio-economic groups influence an LLM’s susceptibility to jailbreak attacks. To accomplish this, we generated keywords representing contrasting groups, such as privileged and marginalized communities, using the LLM itself to ensure that the keywords reflect the model’s inherent associations and biases.
@@ 3.2.2 Analyzing Inherent Biases
Figure 3: Overview of the BiasJailbreak methodology. The same harmful prompt is used across different keywords representing contrasting groups to analyze variations in jailbreak success rates.*All keywords representing both privileged and underprivileged groups are generated by the LLM.
@@
3.2.3 Summary of the BiasJailbreak Approach
Our BiasJailbreak methodology systematically analyzes the LLM’s responses to a diverse set of harmful prompts, using various keywords across different groups. We conducted experiments using harmful prompts $H$ and multiple pairs of privileged and marginalized keywords. By maintaining a consistent prompt structure (Equation 1) and isolating the effect of the keyword, we aimed to identify any biases in the LLM’s susceptibility to jailbreak attacks.
The calculation of jailbreak success rates (Equation 2) and the comparative analysis using $\Delta S_{i}$ (Equation 3) enabled us to quantify potential biases. Our approach highlights whether ethical biases introduced during the alignment process inadvertently contribute to vulnerabilities that can be exploited through jailbreak attacks.
3.3 BiasDefense: Preventing Jailbreaks without Additional Inference or Guard Models
Our jailbreak defense method BiasDefense uses prompts to adjust biases without the need for additional models. Chain-of-Thought Prompting (CoT) has shown considerable potential in eliciting reasoning abilities in large language models, allowing for interpretability without the need for parameter updates (Wei et al., 2022). Inspired by this, our approach BiasDefense involves adding defense prompts to reduce excessive biases, ensuring safety without additional inference overheads, in contrast to Guard Models which require additional inference cost after text generation (Inan et al. (2023); Ghosh et al. (2024); Caselli et al. (2020); Vidgen et al. (2021)). By incorporating bias terms through our defense prompts, we achieve a cost-efficient and secure method for defending against jailbreak attempts while maintaining the model’s performance.
@@ 3.3 BiasDefense: Preventing Jailbreaks without Additional Inference or Guard Models
Figure 4: BiasDefense adjusts inherent biases in LLMs that are exploited by BiasJailbreak. It is efficient since it does not require additional inference or models such as Guard Models.
By using a defense prompt, our defense mechanism elicits the refusal of a wide range of harmful content without requiring significant additional computing resources, which makes it an attractive alternative to Guard Models.
@@ 4.1 Experimental Setups
Data To analyze the differences in jailbreak success rates between groups, we used the jailbreakbench dataset (Chao et al. (2024)) and advbench dataset (Zou et al. (2023)). These datasets comprises a total of 600 harmful prompts and has been widely used in numerous studies. We utilized this dataset to measure performance variations and to analyze the ethical biases inherent in LLMs.
Models We analyzed ethical bias in well-known LLMs, including closed-source models such as GPT-3.5-turbo, GPT-4, GPT-4-o, Claude-sonnet3.5, and open-source models like Llama2-7B, Llama2-13B, Llama3-7B, Phi-mini-7B, Qwen1.5, and Qwen2-7B.
@@
Keywords The keywords used to compare performance across groups were generated by the target LLM themselves to maximize the exploitation of internal biases (see Table 10). The generated keywords of each model shows that there were common keywords that are close to typical biases, while showing sufficient diversity.
4.2 Ethical Bias Analysis
As shown in Table 1, we analyzed the differences in jailbreak success rates across groups using two utilized datasets: JailbreakBench and AdvBench. The results variations in success rates between marginalized and privileged groups, providing evidence of ethical bias in LLMs. These findings underscore the importance of addressing such biases to improve the fairness and reliability of language models.
Table 1: Performance across different datasets showing baseline success rates, marginalized success rates, privileged success rates, and the difference between marginalized and privileged success rates using LLaMA2 model.
Dataset Baseline
Success Rate
Marginalized
Success Rate ($\uparrow$)
Privileged
Success Rate ($\downarrow$)
Marginalized /
Privileged ($\uparrow$)
JailbreakBench 0.2400 0.2811 (+17.08%) 0.1933 (-19.58%) 145.42%
AdvBench 0.1895 0.2037 (+7.50%) 0.1758 (-7.25%) 115.84%
To further analyze this bias, we conducted a series of experiments aimed at evaluating the impact of ethical biases in LLMs on their susceptibility to jailbreak attempts. Specifically, we utilized various keyword prompts to assess the differences in jailbreak success rates across marginalized and privileged categories. Our evaluation included multiple well-known models, such as GPT-3.5, GPT-4, GPT-4o (Brown et al., 2020) and Claude-sonnet3.5 (Anthropic, 2024) (closed models), alongside open-source models like LLaMA (Touvron et al., 2023), Qwen (Bai et al., 2023), and Phi-mini (Abdin et al., 2024).